# Uncovering the Functional Roles of Nonlinearity in Memory

### v1

@@ 4.6 Nonlinear Decoding Enables Compositional Generalization
Finally, we investigate the SCAN benchmark [33], which requires to systematically recombine learned action primitives according to syntactic structures, tapping into executive functions such as manipulation and control of memory content. In this setting, a sequence of linguistic commands (e.g., run opposite left after walk right) must be mapped to a corresponding sequence of low-level actions (e.g., turn right, walk, turn left, turn left, run; Fig. 10). The SCAN grammar is composed of conjunctions (and) and temporal dependencies (after), requiring the model to either execute commands sequentially or invert their order during decoding [33].
@@
We adopt a modular encoder–decoder architecture, assigning a separate AL-RNN to encode the commands and to decode the resulting latent state into action sequences, separately varying the number of nonlinear units $P$ in the encoder and decoder AL-RNNs (out of a total of 128 units each). As for the two previous tasks, purely linear models remain stuck at around $20-30\%$ sequence accuracy (Fig. 7a). Sparse nonlinearity ($P=1,4$) - while improving performance - remains insufficient to solve the task reliably. Performance peaks when both encoder and decoder include moderate nonlinearity ($P=32$), with near-perfect accuracy ($>99.9\%$), and then slightly deteriorates for fully nonlinear models ($P=128,\approx 96\%$).
@@
In addition, we observe a clear asymmetry: a linear encoder can suffice, achieving up to $99\%$ accuracy when paired with a sufficiently nonlinear decoder ($P=64$), while a linear decoder gets stuck below $80\%$ regardless of the encoder. To better understand this behavior, we analyzed the latent structure produced by the encoder. Even when the encoder is purely linear, the final encoded states map onto a large number of distinct subregions in the decoder’s PWL space: for instance, in a model with $P_{\mathrm{enc}}=0$ and $P_{\mathrm{dec}}=64$, the 4,182 test commands map to 4,054 distinct decoder subregions. Thus, while the encoder itself remains linear, end-to-end training organizes the latent representations such that the decoder’s nonlinearity can be exploited effectively.
@@
Second, the combination of a fully nonlinear encoder ($P=128$) and a moderately nonlinear decoder ($P=4$) can also achieve $>99\%$ test accuracy while enabling a clear syntactic interpretation of the latent units. We found that this behavior emerges from the interaction between the expressive encoder, which maps syntactic constructs onto the PWL latent space of the decoder, effectively leveraging its subregions for decoding (Fig. 7). Specifically, a compact subspace of four subregions, characterized by the first two PWL units remaining inactive, captures the majority of intial states corresponding to shorter action sequences (Fig. 7b). In contrast, more complex and lengthier command sequences are consistently mapped to a separate latent subspace, where nonlinear dynamics become more pronounced (Fig. 7c). Within this second subspace, we observe a clear syntactic segregation, with commands containing and and after being mapped to mutually exclusive linear subregions (Fig. 7d). Additionally, we find that during decoding, transitions between composite commands in an action sequence are synchronized with sign flips in the PWL units (Fig. 21).
Collectively, these findings suggest that while symbolic encoding in SCAN can remain largely linear, compositional generalization depends on the PWL structure of the decoder that tiles the latent space into modular, context-sensitive, and syntax-aware subregions. This decomposition reveals a pathway toward interpreting how abstract syntactic operations are organized in more complex sequence tasks.
[…]
@@ 5 Discussion
This work systematically investigates how nonlinearity supports memory and structured computation across a diverse range of sequence modeling tasks. Using AL-RNNs, we identify task-dependent thresholds where sparse nonlinearity dramatically enhances expressivity and memory capacity. Across tasks, simple linear mechanisms form the backbone of memory, while selective nonlinearity enables task and context-specific computations. The PWL structure of the AL-RNN makes these mechanisms transparent, directly linking switches in model dynamics to computation. Our findings challenge the assumption that greater nonlinearity always improves expressivity. We show that fully nonlinear models often perform worse than a combination of linear and nonlinear units. This may partially explain the efficiency of linear SSMs in tasks with long-range dependencies [24, 60, 48], and highlights the benefits of a targeted, sparse approach to nonlinearity in recurrent models.
The perspective that the structured separation of linear and nonlinear dynamics enables the extraction of computational mechanisms is also highly relevant for analyzing the latent structure of neural dynamics [16], which encodes similar mechanisms such as storage, gating, and symbolic manipulation [67, 42, 12]. Our analysis of rodent neural data supports this view: although the task is linearly solvable, jointly modeling behavior and spike activity benefits from nonlinearity to accommodate the variability inherent in empirical neural data. Our results demonstrate how structured models like the AL-RNN can uncover interpretable dynamics in noisy, real-world settings, while also highlighting that much of the complexity in neural data may actually stem from noise, drift, or unobserved inputs, not the task itself.
We focus primarily on memory in RNNs but also demonstrate that operations relevant to reasoning, such as conditional processing and context sensitivity, emerge naturally from the underlying PWL dynamics. Mechanisms like gating and symbolic recombination are enabled by specific nonlinear transitions, illustrating how RNNs can perform structured reasoning without explicit symbolic representations. The modular organization of latent space through PWL units, akin to cortical processing [20, 3], offers a structured pathway for mechanistic interpretability.
@@
Our study covers a structured but limited set of memory tasks. While the identified dynamics capture important memory strategies, many real-world tasks, particularly in language, may blend these mechanisms more intricately. Extending our analysis to more complex architectures, including structured SSMs like Mamba [22] or Transformer-based models [65], represents a promising path forward for bridging interpretability and SOTA sequence models.
@@ Acknowledgements
This research was funded by the German Research Foundation (DFG) within the framework of the Collaborative Research Centre subprojects A06 and B08. We thank the Hector II Foundation for additional support.
@@ A.1.1 Training Details
The primary focus of this study was to investigate how nonlinearity affects performance. To ensure that our observations were not confounded by latent dimensionality, we used relatively high-dimensional AL-RNN models, scanning over several values of the latent size $M$ in the range of 10 to 100 (up to $128$ for SCAN given the higher complexity of the task). This allowed us to operate in a regime where performance was not limited by the latent state capacity. The main hyperparameter of interest was the number of nonlinear units $P$, which was systematically varied across experiments to assess its impact on task performance and dynamical properties. Hyperparameter scans for $P$ are reported directly in the results figures of the main text.
For optimization, we employed standard settings consistent with the original AL-RNN implementation [9], including initialization strategies from the official repository. We initialized the self-connection weights in $\bm{A}$ with small random values to prevent premature convergence to unstable dynamics, while the weights $\bm{W}$, the bias term $\bm{h}$, the input matrix $\bm{C}$ and readout matrix $\bm{D}$ (mapping to logit scores for classification) were sampled from a Gaussian distribution with zero mean and a standard deviation of $0.01$. For optimization, we employed Adam [32] with a learning rate of $0.001$, which was further reduced during training by a cosine annealing learning rate scheduler. We used these settings for all datasets and did not perform fine-grained tuning.
During training, the model parameters were trained using standard backpropagation through time (BPTT). An independent validation set, comprising $10\%$ of the training data, was used for model selection. For datasets with pre-existing validation splits (e.g., sMNIST, Speech Commands, IMDb, SCAN), this was taken from the training set. For datasets without predefined splits, a validation set was artificially created. Other hyperparameters, such as batch size and learning rate scheduler, were set to conventional values that ensured stable convergence across all tasks.
@@
All models presented in this study are relatively lightweight, using single-layer AL-RNNs with a maximum of 128 units. Due to this efficiency, all experiments could be conducted on a single CPU, with total training times for each model not exceeding 12 hours. Hence, the memory footprint remained well within the capacity of standard computing setups (less than $8$GB RAM). The modest computational requirements ensure that our approach is easily replicable on conventional hardware.
@@ A.1.2 Manifold Attractor Regularization
To promote the reconstruction of both fast and slow time scales - and associated memory - in the latent space, we regularize a subset $M_{\mathrm{reg}}\leq M$ of the latent states, as suggested in Schmidt et al. [57], according to:
@@
Here, $\tilde{\bm{A}}$ represents the effective diagonal term for each unit: for linear units, the diagonal elements are derived directly from the $\bm{W}$ matrix, as $\bm{A}$ is zero in this case. For nonlinear units, the diagonal contribution is computed as the sum of their respective entries in $\bm{A}$ and $\bm{W}$. This ensures that the regularization correctly constrains self-connections to be near $1$, irrespective of the unit’s type. The regularization loss is added to the total loss with a scaling factor $\tau$ for each setting, which we scanned for $\tau\in\{0,0.1,0.5,1.0,5.0,10\}$ for the integration tasks (Fig. 8). We found a value of $0.1$ to overall lead to good performance, which we settled on it for all results presented in the paper.
By constraining the self-connections of certain units to be near 1 and suppressing their cross-connections, these sub-units function as near-perfect integrators, capable of encoding arbitrarily slow time scales. Dynamically, this regularization encourages the formation of a continuous manifold of marginally stable fixed points, supporting long-term integration without requiring architectural modifications such as gating. It enables efficient memory retention even in simple RNN architectures, and has been shown to outperform LSTMs and other RNN architectures on long-range tasks [57]. As recommended by Schmidt et al. [57], we applied the regularization to half of the latent units $M$, hence primarily targeting the linear subspace, depending on the number of $P$ PWL units.
Figure 8: Manifold Attractor Regularization (MAR, [57]) leads to a strong performance boost on sMNIST and Speech Commands Task, while not being overly sensitive to the exact value of the regluarization strength $\tau$.
@@ A.1.3 Encoders for Text, Audio, and Images
For the IMDB sentiment classification task, we use pre-trained $100D$ GloVe embeddings [51] to initialize word vectors, followed by a multi-layer perceptron (MLP) as a nonlinear encoder. This additional nonlinear mapping allows the model to capture high-level semantic abstractions from raw word embeddings, which is crucial for sentiment classification [31, 69]. Without this structured transformation, the model struggles to resolve complex syntactic and semantic relationships in the text.
@@
For the sMNIST task, we apply a series of stacked local 1D convolutional layers to the pixel sequences. These 1D convolutions respect the sequential nature of the input while extracting local stroke-based features that are beneficial for digit recognition.
@@
For the Speech Commands dataset, we transformed the raw audio signals into Mel Frequency Cepstral Coefficients (MFCCs), a widely used feature representation for speech processing [14]. MFCCs provide a compact and noise-robust representation, emphasizing the perceptually relevant aspects of speech. This transformation also downsamples the sequence length based on the MFCC hyperparameters. In our setup, we select parameters that preserve moderate sequence lengths comparable to sMNIST $16k$ time steps. The transformed MFCC features are then passed through a series of stacked local 1D convolutional layers to extract temporal patterns. To ensure consistency across samples, the MFCC features are normalized using global mean and standard deviation, computed only over the training data.

### v2

@@ 4.6 Nonlinear Decoding Enables Compositional Generalization
Finally, we investigate the SCAN benchmark (Lake and Baroni, 2018), which requires to systematically recombine learned action primitives according to syntactic structures, tapping into executive functions such as memory manipulation and control. In this setting, a sequence of linguistic commands (e.g., run opposite left after walk right) must be mapped to a corresponding sequence of low-level actions (e.g., turn right, walk, turn left, turn left, run; Fig. 8). The SCAN grammar is composed of conjunctions (and) and temporal dependencies (after), requiring the model to either execute commands sequentially or invert their order during decoding (Lake and Baroni, 2018).
Figure 6: a: Sequence-level accuracy (in $\%$) vs. number of nonlinear units $P$ in encoder (x-axis) and decoder (y-axis). b: Bitcode distribution of initial states of the decoder indicate a dominant subspace for which the first two units are zero. c: Average action sequence length associated with initial states lying in different subregions. d: Relative frequency of commands containing and and after within subregions shows clear syntactical separation.
@@
We adopt a modular encoder–decoder architecture (illustrated in Fig. 8), assigning two separate AL-RNNs to encode the commands and to decode the resulting latent state into action sequences, separately varying the number of nonlinear units $P$ in the encoder and decoder AL-RNNs (out of a total of 128 units each). Purely linear models remain stuck at around $20-30\%$ sequence accuracy (Fig. 6a). Sparse nonlinearity ($P=1,4$) - while improving performance - remains insufficient to solve the task reliably. Performance peaks when both encoder and decoder include moderate nonlinearity ($P=32$), with near-perfect accuracy ($99.9\%$), and then slightly deteriorates for fully nonlinear models ($P=128,\approx 96\%$). In addition, we observe a clear asymmetry: a linear encoder can suffice, achieving up to $99\%$ accuracy when paired with a sufficiently nonlinear decoder ($P=64$), while a linear decoder remains below $80\%$ regardless of the encoder.
@@
To better understand this behavior, we analyzed the latent structure produced by the encoder. Even when the encoder is purely linear, the final encoded states map onto a large number of distinct subregions in the decoder’s PWL space: thus, while the encoder itself remains linear, end-to-end training organizes the latent representations such that the decoder’s nonlinearity can be exploited effectively.
@@
Second, the combination of a fully nonlinear encoder ($P=128$) and a moderately nonlinear decoder ($P=4$) can also achieve $99\%$ test accuracy while enabling a clear syntactic interpretation of the latent units. We found that this behavior emerges from the interaction between the expressive encoder, which maps syntactic constructs onto the PWL latent space of the decoder, effectively leveraging its subregions for decoding (Fig. 6). Specifically, a compact subspace of four subregions, with the first two PWL units inactive, captures most initial states for shorter sequences (Fig. 6b). In contrast, more complex and lengthier command sequences are consistently mapped to a separate latent subspace, where nonlinear dynamics become more pronounced (Fig. 6c). Within this second subspace, we observe a clear syntactic segregation, with commands containing and and after being mapped to mutually exclusive linear subregions (Fig. 6d). Additionally, we find that during decoding, transitions between composite commands in an action sequence are synchronized with sign flips in the PWL units (Fig. 26). Collectively, these findings suggest that while symbolic encoding in SCAN can remain largely linear, compositional generalization depends on the decoder’s PWL structure tiling latent space into modular, syntax-aware subregion. This decomposition reveals a pathway toward interpreting how abstract syntactic operations are organized in more complex sequence tasks.
Table 1: Recurrent motifs identified by AL-RNNs across tasks, linking dynamics to computation.
Task Motif(s) Computation
IMDb Slow mode Sentiment integration
sMNIST, SpeechCmd Class-dependent slow modes Evidence accumulation & pattern-specific routing
Addition Slow mode + gating Selective integration of marked inputs
Copy (fixed delay) Limit cycle + subregion transition Stable storage & timed recall
Copy (variable delay) Limit cycle + subregion transitions Flexible storage & sequence decoding
Context-dependent integration Rule-specific slow modes Rule-based integration
CRCNS-PFC-1 (rodent data) Rule-specific slow modes Rule-based routing
SCAN Syntax-dependent subregion partitioning Syntax-based modular decoding
Multi-task training Task-specific subregion allocation Compositional reuse of nonlinear motifs
@@ 5 Discussion
This work asks whether—and how—systematically attenuating nonlinearity can make the computational mechanisms underlying sequence learning tasks explicit and tractable. Instead of treating nonlinearity as a binary design choice, we study how progressively tuning its strength reveals which aspects of temporal computation genuinely require nonlinear recurrence, and which can be supported by predominantly linear dynamics.
We pursue this question using the AL-RNN, whose piecewise-linear structure allows nonlinearity to be precisely controlled and directly observed. Varying the number of nonlinear units enables a step-wise interpolation between linear and nonlinear recurrence, while the induced partitioning of state space enables bitcode-based analyses and analytical characterization of the underlying dynamics. This combination allows us to directly identify mechanisms—such as gating, switching, and regime changes—through which nonlinear recurrence contributes to sequence computation.
Across tasks, we observe a consistent organizational principle, summarized in Table 1: memory itself is often implemented via a slow linear mode, while computational operations are layered on top through sparse nonlinear mechanisms, often realized as discrete switches/gates. The PWL structure of the AL-RNN makes these mechanisms transparent, directly linking switches in model dynamics to computation. Our multi-task experiments further reveal that sparsely nonlinear models enable sample-efficient transfer learning, with computational reuse across tasks directly linked to shared linear subregions. This suggests that multi-task settings provide a promising avenue for understanding how neural systems discover and share computational primitives across related problems.
Beyond reducing interpretability, we find that for some tasks, fully nonlinear models indeed perform worse than a mix of linear and nonlinear units, which we could link to less stable optimization. We interpret this as a useful inductive bias of sparsely nonlinear models, resting on converging observations: low-$P$ models tend to yield compact, interpretable solutions, whereas fully nonlinear models fragment computation across many subregions; fully nonlinear models often end up “linearizing” units during training, which is harder to optimize than starting with linear units directly; and these effects are robust across activation functions. While tailored regularization or initialization could in principle mitigate these gaps, our results provide mechanistic insight into why sparse nonlinearity can perform better “out of the box”, and suggest that the sparse introduction of nonlinearity in otherwise linear backbones could be leveraged as a design principle for more sophisticated sequence modeling architectures.
A natural concern is that our findings may be overly specific to the AL-RNN and its PWL activations. We use the AL-RNN in a deliberately reductionist way - not to argue that PWL nonlinearities are the “right” way to solve these tasks, but to break the model down into the simplest components needed for understanding. The PWL formulation makes it easy to see when and how the network’s dynamics change, which in turn makes the underlying computations more transparent. Our goal is therefore not necessarily invariance at the level of low-level mechanisms, since different nonlinearities can realize the same computation in different ways, but invariance at the level of functional roles. From this perspective, the patterns we observe—largely linear memory dynamics interrupted by occasional nonlinear events that gate, switch, or change regimes—are not specific to the AL-RNN, but reflect common strategies that recur across tasks and activation functions.
More broadly, this structured separation of linear and nonlinear dynamics provides a useful lens for analyzing latent neural dynamics where similar mechanisms underlie storage, gating, and context-dependent computation (Durstewitz et al., 2023, Wang, 2008, Mante et al., 2013, Cho et al., 2014). Our results demonstrate how structured models can expose such mechanisms even in noisy, real-world settings, and suggest that modular, sparsely nonlinear latent representations - reminiscent of cortical organization (Fusi et al., 2016, Goyal et al., 2021) - offer a promising pathway toward mechanistic interpretability in both artificial and biological systems. More broadly, the structured separation of linear and nonlinear dynamics provides a useful lens for analyzing latent neural dynamics where similar mechanisms underlie storage, gating, and context-dependent computation (Durstewitz et al., 2023, Wang, 2008, Mante et al., 2013, Cho et al., 2014). Our results demonstrate how structured models can expose such mechanisms even in noisy, real-world settings, and suggest that modular, sparsely nonlinear latent representations - reminiscent of cortical organization (Fusi et al., 2016, Goyal et al., 2021) - offer a promising pathway toward mechanistic interpretability in both artificial and biological systems, suggesting potential applications to understanding how language models perform complex reasoning.
@@
Our study covers a structured but limited set of sequence tasks. While the identified dynamics capture important computational strategies, many real-world tasks, particularly in language, may blend these mechanisms more intricately. While we conducted preliminary analyses of sparse nonlinearity in more complex architectures, including S4 (Gu et al., 2022) and linear Transformers (Katharopoulos et al., 2020), understanding the role of nonlinearity in these models presents distinct challenges, since here, linear recurrence and linear attention mechanisms are integral to optimization and cannot be straightforwardly replaced with AL-RNN-style sparse PWL reccurence. Moreover, when nonlinearity is introduced through layer-wise transformations rather than within the recurrence itself, the connection to interpretable dynamical objects (fixed points, eigenspectra, flow fields) becomes less clear. Although we observed benefits of strategically placed sparse nonlinearity even in these architectures for the tasks tested (Fig. 27), a comprehensive mechanistic understanding of these effects requires future work. Extending our interpretability framework to modern architectures represents a promising path toward bridging mechanistic understanding with state-of-the-art sequence models.
@@ Acknowledgements
This research was funded by the German Research Foundation (DFG) within the Collaborative Research Centre 265 subprojects A06 and B08, the Federal Ministry of Research, Technology and Space (BMFTR), the Baden-Württemberg Ministry of Science as part of the Excellence Strategy of the German Federal and State Governments, and the DYNAMIC Centre, funded by the LOEWE program of the Hessian Ministry of Science and Arts (Grant Number: LOEWE 1/16/519/03/09.001(0009)/98).
Reproducibility
Code for experiments and datasets is publicly available under https://github.com/ManuelBrenner/Linear_Nonlinear_Memory.
@@ Appendix A Appendix
This appendix provides further methodological details and extended results. Sect. A.1 describes training procedures, regularization, encoders, and spike data modeling. Sect. A.2 covers analytical methods for bitcode analysis, stability, Lyapunov exponents, and flow fields. Sect. A.3 details all datasets. Sect. A.4 presents extended results across all tasks, including analysis of a stimulus selection task on rodent spike data (A.4.5).
@@ A.1.1 Training Details
The primary focus of this study was to investigate how nonlinearity affects performance. To ensure that our observations were not confounded by latent dimensionality, we used relatively high-dimensional AL-RNN models, scanning over several values of the latent size $M$ in the range of 10 to 100 (up to $128$ for SCAN given the higher complexity of the task). This allowed us to operate in a regime where performance was not limited by the latent state capacity, and where a range of values for $P$ could be tested. We also conducted targeted sweeps over the manifold-attractor regularization strength $\tau\in\{0,0.01,0.1,0.5,1.0,5.10\}$ (see Sect. A.1.2). In practice, $\tau=0.1$ consistently yielded good results, and we therefore fixed this value across tasks to reflect that our results are not dependent on overly specific hyperparameter tuning. Importantly, the qualitative performance patterns across different levels of nonlinearity were robust to variations in both $M$ and $\tau$. The main hyperparameter of interest remained the number of nonlinear units $P$, which was systematically varied across experiments to assess its impact on task performance and dynamical properties. Hyperparameter scans for $P$ are reported directly in the results figures of the main text.
For optimization, we employed standard settings consistent with the original AL-RNN implementation (Brenner et al., 2024a), including initialization strategies from the official repository. We initialized the self-connection weights in $\bm{A}$ with small random values to prevent premature convergence to unstable dynamics, while the weights $\bm{W}$, the bias term $\bm{h}$, the input matrix $\bm{C}$ and readout matrix $\bm{D}$ (mapping to logit scores for classification) were sampled from a Gaussian distribution with zero mean and a standard deviation of $0.01$. For optimization, we employed Adam (Kingma and Ba, 2017) with a learning rate of $0.001$, which was further reduced during training by a cosine annealing learning rate scheduler. We used these settings for all datasets and did not perform fine-grained tuning.
During training, the model parameters were trained using standard backpropagation through time (BPTT). An independent validation set, comprising $10\%$ of the training data, was used for model selection. For datasets with pre-existing train/test splits (e.g., sMNIST, Speech Commands, IMDb, SCAN), this was taken from the training set. For datasets without predefined splits, a validation set was artificially created. Other hyperparameters, such as batch size and learning rate scheduler, were set to conventional values that ensured stable convergence across all tasks.
@@
All models presented in this study are relatively lightweight, using single-layer AL-RNNs with a maximum of 128 units. Due to this efficiency, all experiments could be conducted on a single CPU, with total training times for each model not exceeding 12 hours. Hence, the memory footprint remained well within the capacity of standard computing setups (less than 88GB RAM). The modest computational requirements ensure that our approach is easily replicable on conventional hardware.
@@ A.1.2 Manifold Attractor Regularization
To promote the reconstruction of both fast and slow time scales - and associated memory - in the latent space, we regularize a subset $M_{\mathrm{reg}}\leq M$ of the latent states, as suggested in Schmidt et al. (2021), according to:
@@
Here, $\tilde{\bm{A}}$ represents the matrix with the effective diagonal term for each unit: for linear units, the diagonal elements are derived directly from the $\bm{W}$ matrix, as $\bm{A}$ is zero in this case. For nonlinear units, the diagonal contribution is computed as the sum of their respective entries in $\bm{A}$ and $\bm{W}$. This ensures that the regularization correctly constrains self-connections to be near $1$, irrespective of the unit’s type. The regularization loss is added to the total loss with a scaling factor $\tau$ for each setting.
By constraining the self-connections of certain units to be near 1 and suppressing their cross-connections, these sub-units function as near-perfect integrators, capable of encoding arbitrarily slow time scales. Dynamically, this regularization encourages the formation of a continuous manifold of marginally stable fixed points, supporting long-term integration without requiring architectural modifications such as gating. It enables efficient memory retention even in simple RNN architectures, and has been shown to outperform LSTMs and other RNN architectures on long-range tasks (Schmidt et al., 2021). As recommended by Schmidt et al. (2021), we applied the regularization to half of the latent units $M$, hence primarily targeting the linear subspace.
Figure 7: Manifold Attractor Regularization (MAR, Schmidt et al. (2021)) leads to a strong performance boost on sMNIST and Speech Commands Task, while not being overly sensitive to the exact value of the regluarization strength $\tau$.
@@ A.1.3 Encoders for Text, Audio, and Images
For the IMDB sentiment classification task, we used pre-trained $100D$ GloVe embeddings (Pennington et al., 2014) to initialize word vectors, followed by a multi-layer perceptron (MLP) as a nonlinear encoder. This additional nonlinear mapping allows the model to capture high-level semantic abstractions from raw word embeddings that aid with sentiment classification (Kim, 2014, Zhou et al., 2015) and enable the AL-RNN to focus on evidence integration, leading to a dynamically simple linear process.
@@
For the sMNIST task, we applied a series of stacked local 1D convolutional layers to the pixel sequences. These 1D convolutions respect the sequential nature of the input while extracting local stroke-based features that are beneficial for digit recognition.
@@
For the Speech Commands dataset, we transformed the raw audio signals into Mel Frequency Cepstral Coefficients (MFCCs), a widely used feature representation for speech processing (Davis and Mermelstein, 1980). MFCCs provide a compact and noise-robust representation, emphasizing the perceptually relevant aspects of speech. This transformation also downsamples the sequence length based on the MFCC hyperparameters. In our setup, we select parameters that preserve moderate sequence lengths comparable to sMNIST (784 time steps). The transformed MFCC features are then passed through a series of stacked local 1D convolutional layers to extract temporal patterns. To ensure consistency across samples, the MFCC features are normalized using global mean and standard deviation, computed only over the training data.